[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Oct 05, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Oct 6 03:23:22 UTC 2016


HSU Shift Log for Wed Oct 05 2016 23:23:20 GMT-0400 (EDT)

Shift: R

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: None.

Other Techincal issues: Scatterometer data out since 2315Z in nmap2
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161006/8d915d66/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list